# Learning Strategy Representation for Imitation Learning in Multi-Agent Games

### v1

@@ Conclusion
In this work, we proposed an effective framework, STRIL, to extract the representations of the offline trajectories and enhance imitation learning methods in multi-agent games. We designed a P-VRNN network, which shows extraordinary results in learning the strategy representations of trajectories without requiring player identification. We then defined two indicators, RI and EL, for imitation learning. We can estimate RI and EL by utilizing the strategy representation and subsequently filter the offline dataset with the indicators. The imitation learning algorithms show significant performance improvements with the filtered datasets. The limitations and broader impacts are provided in the Appendix.

### v2

@@ Conclusion
In this work, we proposed an effective framework, STRIL, to extract the representations of the offline trajectories and enhance imitation learning methods in multi-agent games. We designed a P-VRNN network, which shows extraordinary results in learning the strategy representations of trajectories without requiring player identification. We then defined two indicators, RI and EL, for imitation learning. We can estimate RI and EL by utilizing the strategy representation and subsequently filter the offline dataset with the indicators. The imitation learning algorithms show significant performance improvements with the filtered datasets.
@@
Acknowledgments
This work was supported in part by the National Natural Science Foundation of China under Grants 72293575, as well as by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No. RS-2024-00410082)